Find the x intercepts.
y=x^2-9x+20

Answer by THANApHD(104)

intercepts mean th x's values when the y = 0.

0=x^2-9x+20
0=x^2-5x-4x+20
0=x(x-5)-4(x-5)
0=(x-4)(x-5)
so either one this must be zero
x-4 = 0 or x-5=0
x= 4 or x=5